write the slope-intercept equation for the line that passes through (5,4) and (7,8) please show all of your work
Answer by mananth(16946)

x1 y1 x2 y2
5 4 7 8

slope m = (y2-y1)/(x2-x1)
(8-4)/(7-5)
( 4 / 2 )
m= 2

Plug value of the slope and point (5,4) in
Y = m x + b
4.00 = 10 + b
b= 4 - 10
b= -6
So the equation will be
Y=2x-6
m.ananth@hotmail.ca